# Modeling Advection-Dominated Flows with Space-Local Reduced-Order Models

### v2

@@ 6 Conclusions
In this work, we presented a novel way of constructing ROMs for PDEs describing advection-dominated problems. The key properties of our proposed ROM are: sparsity and generalizability through a space-local ROM with overlapping subdomains; stability by embedding energy conservation in the ROM. The space-local ROMs are achieved by building a POD basis within each subdomain, as compared to a single common basis for the standard space-global approach. In [25] it was shown that such a basis results in a sparse and computationally efficient ROM. We modified this approach to generate a common local basis for the entire domain, similar to a finite element basis. By generating a common basis we achieve generalizability, as in this way a feature observed in one subdomain can now be represented in any of the subdomains. In addition, we introduced overlapping subdomains to prevent discontinuities at the subdomain boundaries. To test our methodologies we made use of the linear advection equation in both 1D and 2D. One of the properties of this system is that the energy is conserved. Our space-local ROMs satisfy this property exactly, even when the basis is non-orthogonal (as is the case for the overlapping approach).
@@
We observed that the resulting space-local ROMs generalize much better for advection-dominated problems than the standard space-global approach. We also showed that such a space-local approach produces sparser and, thus, more computationally efficient ROMs. In addition, we showed that the space-local ROMs also satisfy energy conservation and allow for larger time steps. The latter further decreases the computational cost of the ROMs. Regarding the computational time, we observed an improvement between one and two orders of magnitude with respect to the FOM in the 2D test case. By introducing overlapping subdomains we show that we obtain smoother approximations of the solution and require a fewer number of basis functions to represent the solution compared to the non-overlapping approach. However, this comes at the cost of solving a linear system at each time step. Depending on the application, either overlapping or non-overlapping subdomains might be most suitable. If smoothness of the solution is required overlapping subdomains might be more suitable, while if computational efficiency is of higher priority non-overlapping subdomains are likely preferred.
@@
For future research, we consider the extension of this methodology to an actual turbulence test case described by the Navier-Stokes equations. To achieve energy-conservation one can make use of the observation presented in [9], where it was shown that carrying out a Galerkin projection on an energy-conserving FOM (with quadratic non-linearity) resulted in an energy-conserving ROM. However, one important requirement is that the POD basis needs to be divergence free, meaning that the space-local POD basis also needs to be divergence free. Currently, this is still being actively researched in our group. In addition, one could consider reusing the learned space-local basis for different spatial domains than represented in the training data. Another possible research direction would be to circumvent or speed up the solution of the linear system required in LO-POD. For this purpose, one could possibly take inspiration from the finite element community [24].
CRediT authorship contribution
Data availability
The code used to generate the training data, construct the ROMs, and replicate the presented experiments can found at https://github.com/tobyvg/local_POD_overlap.jl.
Acknowledgements
This publication is part of the project “Unraveling Neural Networks with Structure-Preserving Computing” (with project number OCENW.GROOT.2019.044 of the research programme NWO XL which is financed by the Dutch Research Council (NWO)). Part of this publication is funded by Eindhoven University of Technology. Finally, we thank the reviewers for their feedback, enhancing the quality of the article.
During the preparation of this work the author(s) used ChatGPT in order to improve language and grammar. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@ Acknowledgements
global proper orthogonal decomposition
@@
Here we discuss the kernel that is used to construct the LO-POD basis for the 2D advection test case. This kernel is required to obtain basis functions that smoothly decay to zero at the edge of the subdomain. To achieve this we use a standard bump function:
@@
to construct the kernel [38]. This function, along with its derivatives, smoothly decays to zero as $|x|$ approaches $1$. Next, we apply normalization
@@
where $\mathbf{x}=(x,y)^{T}\in\mathbb{R}^{2}$ are the spatial coordinates. For the presentation of this kernel we exploit the fact that we use a uniform grid. This allows use to conveniently subdivide the domain into square subdomains $\Omega_{i}$ of the same size. Let $\boldsymbol{\alpha}^{i},\boldsymbol{\beta}^{i}\in\mathbb{R}^{2}$ denote the coordinates of the bottom-left and top-right vertex. Each subdomain overlaps with its 4 neighboring subdomains which share a vertex in the middle of the considered domain. The kernel associated with subdomain $\Omega_{i}$ is
@@ Appendix B Choice of subdomains and modes for 2D test case
In this section we present the projection error on the training and validation set for the 2D advection test case. This is presented in Figure 15.
@@
Based on these result we select the number of modes $q$ which first surpasses the projection error of $10^{-2}$ for $I=8\times 8$. This value of $I$ is chosen as it generalizes well, i.e. the performance on the validation set is similar to the training set and the resulting ROM is sparser than for $I=4\times 4$. For L-POD this corresponds to $q=20$ and for LO-POD to $q=15$.
@@ Appendix C Time step size
In Figure 16 we present the error of ROMs integrated using different time step sizes. The maximum time step sizes are selected so that performance degradation or occurrence of instabilities is avoided. These are presented in Table 1.
Figure 16: Solution error averaged over the simulation for the 2D advection test case the ROMs for each time step size. Results for a finite volume (FV) discretization on a coarser grid than the FOM resolution ($256\times 256$) are also depicted. Absence of markers indicates an unstable simulation. The LO-POD ROM is integrated using a Crank-Nicolson scheme, as opposed to RK4 for the others, and is therefore unconditionally stable.

### v3

@@ 6 Conclusions
In this work, we presented a novel way of constructing ROMs for PDEs describing advection-dominated problems. The key properties of our proposed ROM are: sparsity and generalizability through a space-local ROM with overlapping subdomains; stability by embedding energy conservation in the ROM. The space-local ROMs are achieved by building a POD basis within each subdomain, as compared to a single common basis for the standard space-global approach. In [farhat_local_POD], it was shown that such a basis results in a sparse and computationally efficient ROMs. As suggested in [CHUNG2024DDROM], we modified this approach to generate a common local basis for the entire domain, similar to a finite element basis. This work differs from [CHUNG2024DDROM] in that we applied this to time-dependent PDEs, as opposed to steady-state problems. In addition, we introduced overlapping subdomains to ensure a smooth representation of the solution within the space-local POD basis. By generating a common basis, we achieve generalizability in time, as in this way, a feature observed in one subdomain can now be represented in any of the subdomains. In addition, we introduced overlapping subdomains to prevent discontinuities at the subdomain boundaries. To test our methodologies, we made use of the linear advection equation in both 1D and 2D. One of the properties of this system is that the energy is conserved. Our space-local ROMs satisfy this property exactly, even when the basis is non-orthogonal (as is the case for the overlapping approach).
@@
We observed that the resulting space-local ROMs generalize much better for advection-dominated problems than the standard space-global approach. We also demonstrated that such a space-local approach yields sparser and, consequently, more computationally efficient ROMs. In addition, we showed that the space-local ROMs also satisfy energy conservation and allow for larger time steps. The latter further decreases the computational cost of the ROMs. Regarding computational time, we observed an improvement of one to two orders of magnitude with respect to the FOM in the 2D test case. By introducing overlapping subdomains, we demonstrated that we obtain smoother approximations of the solution and require fewer basis functions to represent it compared to the non-overlapping approach. However, this comes at the cost of solving a linear system at each time step. Depending on the application, either overlapping or non-overlapping subdomains might be most suitable. If smoothness of the solution is required, overlapping subdomains might be more suitable, while if computational efficiency is of higher priority, non-overlapping subdomains are likely preferred.
To further evaluate the generalizability of the space-local POD approaches, we applied it to the 2D incompressible Navier–Stokes equations in a Kolmogorov flow setup. In this case we solely considered the ability of POD basis to represent the solution, without building the ROMs. We consider constructing the actual ROM for this non-linear 2D test case outside the scope of this research, as it requires dealing with the divergence freeness condition, and possibly introducing hyperreduction methods for the non-linear term [hyperreduction_1, hyperreduction_2]. Regarding the representation of the solution space-local methods (L-POD and LO-POD) clearly outperformed G-POD, which showed strong overfitting to the training data. In contrast, the local variants maintained low projection errors and accurately reproduced flow features on unseen, temporally decorrelated data. These results indicate that the space-local framework generalizes well for advection-dominated systems and holds strong potential for reduced-order modeling of turbulent flows.
@@
For future research, we consider constructing a space-local ROM for an actual turbulence test case described by the Navier-Stokes equations. To achieve energy conservation, one can make use of the observation presented in [podbenjamin], where it was shown that carrying out a Galerkin projection on an energy-conserving FOM (with quadratic non-linearity) resulted in an energy-conserving ROM. However, one important requirement is that the POD basis needs to be divergence-free, meaning that the space-local POD basis also needs to be divergence-free. Currently, this is still being actively researched in our group. To resolve this, we could take inspiration from [CHUNG2024DDROM] and construct a POD basis, not only for the velocity, but also for the pressure [rosenberger2022pressureenergyconsistentromsincompressible]. Another possible research direction would be to circumvent or speed up the solution of the linear system required in LO-POD. For this purpose, one could possibly take inspiration from the finite element community [FEM_book]. Finally, the application of the space-local POD methods to steady-state problems could be investigated. The work in [CHUNG2024DDROM] offers an interesting starting point for this, as they investigated the generalization of the space-local POD basis when increasing the size of the domain for steady-state problems. An interesting research avenue would be to compare the space-local ROMs to space-global ones, when the spatial domain is kept the same and one of the model parameters is varied.
@@
space-global proper orthogonal decomposition
@@
CRediT authorship contribution
Data availability
The code used to generate the training data, construct the ROMs, and replicate the presented experiments can be found at https://github.com/tobyvg/local_POD_overlap.jl.
Acknowledgements
This publication is part of the project “Unraveling Neural Networks with Structure-Preserving Computing” (with project number OCENW.GROOT.2019.044 of the research programme NWO XL, which is financed by the Dutch Research Council (NWO)). Part of this publication is funded by Eindhoven University of Technology. Finally, we thank the reviewer and editor for their feedback, enhancing the quality of the article.
During the preparation of this work, the author(s) used ChatGPT to improve language and grammar. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@ Acknowledgements
Here we discuss the kernel used to construct the LO-POD basis for the 2D advection test case. This kernel is required to obtain basis functions that smoothly decay to zero at the edge of the subdomain. To achieve this, we use a standard bump function:
@@
to construct the kernel [lee2012introduction_bump]. This function, along with its derivatives, smoothly decays to zero as $|x|$ approaches $1$. Next, we apply normalization
@@
where $\mathbf{x}=(x,y)^{T}\in\mathbb{R}^{2}$ are the spatial coordinates. For the presentation of this kernel, we exploit the fact that we use a uniform grid. This allows us to conveniently subdivide the domain into square subdomains $\Omega_{i}$ of the same size. Let $\boldsymbol{\alpha}^{i},\boldsymbol{\beta}^{i}\in\mathbb{R}^{2}$ denote the coordinates of the bottom-left and top-right vertex. Each subdomain overlaps with its four neighboring subdomains, which share a vertex in the middle of the considered domain. The kernel associated with subdomain $\Omega_{i}$ is
@@ Appendix B Choice of subdomains and modes for 2D test case
In this section, we present the projection error on the training and validation sets for the 2D advection test case. This is presented in Figure 18.
@@
Based on these results, we select the number of modes $q$ which first surpasses the projection error of $10^{-2}$ for $I=8\times 8$. This value of $I$ is chosen as it generalizes well, i.e., the performance on the validation set is similar to the training set and the resulting ROM is sparser than for $I=4\times 4$. For L-POD this corresponds to $q=20$ and for LO-POD to $q=15$.
@@ Appendix C Time step size
In Figure 19, we present the error of ROMs integrated using different time step sizes. The maximum time step sizes are selected so that performance degradation or the occurrence of instabilities is avoided. These are presented in Table 1.
Figure 19: Solution error averaged over the simulation for the 2D advection test case, the ROMs for each time step size. Results for a finite volume (FV) discretization on a coarser grid than the FOM resolution ($256\times 256$) are also depicted. Absence of markers indicates an unstable simulation. The LO-POD ROM is integrated using a Crank-Nicolson scheme, as opposed to RK4 for the others, and is therefore unconditionally stable.